mary is 15 years older than herb sister jane six years ago mary was 6 times as old as jane what are their present ages
Answer by kapilsinghi(68)

mary is 15 years older than herb sister jane six years ago mary was 6 times as old as jane what are their present ages
Solution:
let Present age of Jane = x
so Present age of Mary = x+15
6 Years ago
Jane age was x-6
mary Age was x+15 - 6=x+9
according to conditions "six years ago mary was 6 times as old as jane"
x+9 = (x-6)*6
x+9 = 6x - 36
9+36 = 6x - x
45 = 5x
x = 45/5
x = 9
present age of jane = 9
age of mary = 24
kapilsinghi123@gmail.com